Session: Mar 2008 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Tue Mar 18 22:54:05 2008) Success :
No known problems

Medicina (A. Cattani at Wed Mar 19 05:11:01 2008) Success :

Torun (Kaz Borkowski at Wed Mar 19 07:06:26 2008) Success :
Not sure, but it seems prior to this experiment the fringe phase jumps seen here as breaks in the LO lock have been eliminated or considerably reduced in frequency of occurrence. Earlier, Monday late afternoon, the required part of azimuth drive has been procured and the broken one replaced.

Effelsberg (Uwe Bach at Wed Mar 19 09:05:20 2008) Minor failures : HIGH
The 6cm receiver got warm during the day, but the cooling was started again in late afternoon. Tsys dropped from 70K to its nominal value of 30K during the first 3 hours of the observations. On hour (2-3 UT) was lost because of snow.

Onsala (Michael Lindqvist at Wed Mar 19 09:36:58 2008) Success :
No known problems.

Jodrell_Bank (Alastair Gunn at Wed Mar 19 15:14:04 2008) Success :
The MkII telescope was used (Jb2). No known problems.

Noto (P. Cassaro at Thu Mar 27 11:23:36 2008) Success :
No known problems